# Deep Learning for Multivariate Time Series Imputation: A Survey

### v1

@@ 5 Time Series Imputation Toolkits
On the time series imputation task, there are existing libraries providing naive processing ways, statistical methods, machine learning imputation algorithms, and deep learning imputation neural networks for convenient usage.
@@
imputeTS Moritz and Bartz-Beielstein , a library in R provides several naive approaches (e.g., mean values, last observation carried forward, etc.) and commonly-used imputation algorithms (e.g., linear interpolation, Kalman smoothing, and weighted moving average) but only for univariate time series. Another well-known R package, mice Van Buuren and Groothuis-Oudshoorn (2011), implements the method called multivariate imputation by chained equations to tackle missingness in data. Although it is not for time series specifically, it is widely used in practice for multivariate time-series imputation, especially in the field of statistics. Impyute [^2] https://github.com/eltonlaw/impyute and Autoimpute [^3] https://github.com/kearnz/autoimpute both offer naive imputation methods for cross-sectional data and time-series data. Impyute is only with simple approaches like the moving average window, and Autoimpute integrates parametric methods, for example, polynomial interpolation and spline interpolation. More recently, GluonTS Alexandrov et al. (2020), a generative machine-learning package for time series, provides some naive ways, such as dummy value imputation and casual mean value imputation, to handle missing values. In addition to simple and non-parametric methods, Sktime Löning et al. (2019) implements one more option that allows users to leverage integrated machine learning imputation algorithms to fit and predict missing values in the given data, though this works in a univariate way. When it comes to deep learning imputation, PyPOTS Du (2023) is a toolbox focusing on modeling partially-observed time series end-to-end. It contains more than a dozen deep-learning neural networks for tasks on incomplete time series, including eight imputation models so far.
6 Experimental Evaluation and Discussion
In this section, empirical experiments are conducted to evaluate and analyze deep multivariate time series imputation methods from different categories. The results are obtained with a machine with AMD EPYC 7543 32-Core CPU and an NVIDIA GeForce RTX 4090 GPU. All code, including the data preprocessing scripts, model configurations, and training scripts, are publicly available in the GitHub repository https://github.com/WenjieDu/Awesome_Imputation.
6.1 Datasets and Imputation Methods
Specifically, three naive imputation approaches and eight deep-learning neural networks are tested on three real-world datasets (Air Zhang et al. (2017), PhysioNet2012 Silva et al. (2012), and ETTm1 Zhou et al. (2021) in Table 2) which are commonly used in the literature.
Regarding the imputation methods, apart from three naive ways Mean, Median, and LOCF (last observation carried forward) as baselines, the eight following representative deep-learning models are selected from different categories for experimental studies: M-RNN Yoon et al. (2019), GP-VAE Fortuin et al. (2020), BRITS Cao et al. (2018), USGAN Miao et al. (2021), CSDI Tashiro et al. (2021), TimesNet Wu et al. (2023a), Transformer Du et al. (2023), and SAITS Du et al. (2023). Experiments are performed with PyPOTS [^4] https://pypots.com Du (2023) and all the above imputation methods are instantly available in the toolbox. Moreover, for a fair comparison, hyperparameters of all deep learning methods are optimized by the tuning functionality in PyPOTS.
Air PhysioNet2012 ETTm1
Number of samples 1,458 11,988 722
Sequence length 24 48 96
Number of features 132 37 7
Original missing rate 1.6% 80.0% 0%
Table 2: The general information of the three preprocessed datasets. Note the detailed descriptions are available in the code repository.
Table 3: The MAE and MSE comparisons between imputation methods on the datasets Air, PhysioNet2012, and ETTm1. The reported values are means $\pm$ standard deviations of five runs.
Method Air PhysioNet2012 ETTm1 MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ Mean 0.692$\pm$0.000 0.970$\pm$0.000 0.702$\pm$0.000 0.954$\pm$0.000 0.663$\pm$0.000 0.809$\pm$0.000 Median 0.660$\pm$0.000 1.027$\pm$0.000 0.685$\pm$0.000 0.991$\pm$0.000 0.657$\pm$0.000 0.825$\pm$0.000 LOCF 0.206$\pm$0.000 0.279$\pm$0.000 0.411$\pm$0.000 0.569$\pm$0.000 0.135$\pm$0.000 0.072$\pm$0.000 M-RNN 0.524$\pm$0.001 0.648$\pm$0.003 0.674$\pm$0.001 0.864$\pm$0.002 0.651$\pm$0.060 1.074$\pm$0.120 GP-VAE 0.280$\pm$0.003 0.266$\pm$0.009 0.400$\pm$0.007 0.433$\pm$0.011 0.290$\pm$0.017 0.178$\pm$0.015 BRITS 0.142$\pm$0.001 0.129$\pm$0.001 0.246$\pm$0.001 0.325$\pm$0.002 0.124$\pm$0.002 0.046$\pm$0.002 USGAN 0.141$\pm$0.001 0.132$\pm$0.001 0.250$\pm$0.001 0.306$\pm$0.001 0.127$\pm$0.005 0.048$\pm$0.003 CSDI 0.105$\pm$0.003 0.153$\pm$0.021 0.211$\pm$0.003 0.260$\pm$0.050 0.157$\pm$0.052 0.292$\pm$0.456 TimesNet 0.159$\pm$0.002 0.172$\pm$0.003 0.266$\pm$0.007 0.272$\pm$0.006 0.113$\pm$0.006 0.027$\pm$0.002 Transformer 0.163$\pm$0.003 0.160$\pm$0.004 0.209$\pm$0.002 0.225$\pm$0.002 0.133$\pm$0.009 0.035$\pm$0.004 SAITS 0.133$\pm$0.002 0.128$\pm$0.001 0.202$\pm$0.002 0.218$\pm$0.002 0.115$\pm$0.011 0.030$\pm$0.006
Method
PR-AUC $\uparrow$ ROC-AUC $\uparrow$
Mean
0.434$\pm$0.016 0.813$\pm$0.009
Median
0.434$\pm$0.018 0.808$\pm$0.014
LOCF
0.425$\pm$0.015 0.804$\pm$0.007
M-RNN
0.424$\pm$0.022 0.807$\pm$0.015
GP-VAE
0.384$\pm$0.018 0.788$\pm$0.008
BRITS
0.428$\pm$0.017 0.821$\pm$0.008
USGAN
0.431$\pm$0.017 0.814$\pm$0.010
CSDI
0.433$\pm$0.017 0.811$\pm$0.005
TimesNet
0.406$\pm$0.012 0.787$\pm$0.013
Transformer
0.446$\pm$0.016 0.807$\pm$0.018
SAITS
0.455$\pm$0.016 0.822$\pm$0.002
Table 3: The MAE and MSE comparisons between imputation methods on the datasets Air, PhysioNet2012, and ETTm1. The reported values are means $\pm$ standard deviations of five runs.
Table 4: The means and standard deviations of classification results in five runs.
6.2 Results and Analysis
Imputation Accuracy Evaluation
Imputation results in error metrics MAE (mean absolute error) and MSE (mean squared error) of twelve methods across three datasets are displayed in Table 4. The numbers tell that the performance of the methods varies on different datasets and there is no clear winner in this study. Further work needs to be done to deeply compare predictive and generative imputation methods. Notably, in cases like the Air and ETTm1 datasets, where data is continuously recorded by sensors and the proportion of missingness is relatively low, the non-parametric LOCF method shows commendable performance. Conversely, in the PhysioNet2012 dataset, which has a high missing rate, deep learning imputation methods markedly outperform statistical approaches. This observation corroborates the capability of deep learning methods to effectively capture complex temporal dynamics and accurately learn data distributions, especially in scenarios with highly sparse, discrete observations.
Downstream Task Evaluation
Generally, the better quality of imputed values represents the better overall dataset quality after imputation. Consequently, in addition to the imputation performance comparison, there is an experiment setting in the literature that evaluates the methods from the perspective of downstream task performance Du et al. (2023). Such a study is adopted in this work as well to help assess the selected methods. A simple LSTM model performs the binary classification task on the PhysioNet2012 dataset where each sample has a label indicating whether the patient in the ICU was deceased. The PhysioNet2012 dataset is processed by the imputation methods and the results are presented in Table 4. PR-AUC (area under the precision-recall curve) and ROC-AUC (area under the receiver operating characteristic curve) are chosen to be the metrics, considering the dataset has imbalanced classes and 14.2% positive samples. Note that the only variable in this experiment is the imputed data.
As shown in Table 4, the classifier can benefit from better imputation on the downstream classification task. The best results from SAITS imputation obtain 5% and 1% gains than the best naive imputation Mean separately on the metrics PR-AUC and ROC-AUC. Please note that such improvements are achieved simply by better imputation, which can be seen as a data-preprocessing step in this experiment. Furthermore, this raises a research question about how to make deep learning imputation models learn from both the imputation task and downstream tasks to obtain a consistent and unified representation from the incomplete time series.
Complexity Analysis
We summarize the time and memory complexity of the deep learning imputation models in Table 5. Additionally, their actual inference time on the test set of PhysioNet2012 is also listed for clear comparison.
Method Computation Memory Running Time
M-RNN $\mathcal{O}(L*K)$ $\mathcal{O}(1)$ 5s
GP-VAE $\mathcal{O}(L*\log{}K)$ $\mathcal{O}(1)$ 1s
BRITS $\mathcal{O}(L)$ $\mathcal{O}(1)$ 9s
USGAN $\mathcal{O}(L)$ $\mathcal{O}(1)$ 9s
CSDI $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ 104s
TimesNet $\mathcal{O}(L*\log{}K)$ $\mathcal{O}(1)$ 1s
Transformer $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ 1s
SAITS $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ 1s
Table 5: Computational and space complexity of imputation models, and their running time in seconds on the PhysioNet2012 test set.
7 Conclusion and Future Direction
This paper presents a systematic review of deep learning models specifically tailored for multivariate time series imputation. We introduce a novel taxonomy to categorize the reviewed methods, providing a comprehensive introduction and an experimental comparison of each. To advance this field, the paper concludes by identifying and discussing the following potential avenues for future research.
@@ 7 Conclusion and Future Direction
Existing imputation algorithms predominantly operate under the MCAR or MAR. However, real-world missing data mechanisms are often more complex, with the MNAR data being prevalent in diverse fields such as IoT devices Li et al. (2023), clinical studies Ibrahim et al. (2012), and meteorology Ruiz et al. (2023). The non-ignorable nature of MNAR indicates a distributional shift exists between observed and true data Kyono et al. (2021). For example, in airflow signal analysis Ruiz et al. (2023), the absence of high-value observations causes MNAR missing mechanism and leads to saturated peaks, visibly skewing the observed data distribution compared to the true underlying one. This scenario illustrates how imputation methods may incur inductive bias in model parameter estimation and underperform in the presence of MNAR. Addressing missing data in MNAR contexts, distinct from MCAR and MAR, calls for innovative methodologies to achieve better performance.
@@
The primary objective of imputing missing values lies in enhancing downstream data analytics, particularly in scenarios with incomplete information. The prevalent approach is the “impute and predicttwo-stage paradigm, where missing value imputation is a part of data preprocessing and followed by task-specific downstream models (e.g. a classifier), either in tandem or sequentially. An alternative method is the “encode and predictend-to-end paradigm, encoding the incomplete data into a proper representation for multitask learning, including imputation and other tasks (e.g. classification and forecasting, etc.). Despite the optimal paradigm for partially-observed time series data still remains an open area for future investigation, the latter end-to-end way turns out to be more promising especially when information embedded in the missing patterns is helpful to the downstream tasks Miyaguchi et al. (2021).
@@
While deep learning imputation algorithms have shown impressive performance, their computational cost often exceeds that of statistical and machine learning based counterparts. In the era of burgeoning digital data, spurred by advancements in communication and IoT devices, we are witnessing an exponential increase in data generation. This surge, accompanied by the prevalence of incomplete datasets, poses significant challenges in training deep models effectively Wu et al. (2023b). Specifically, the high computational demands of existing deep imputation algorithms render them less feasible for large-scale datasets. Consequently, there is a growing need for scalable deep imputation solutions, leveraging parallel and distributed computing techniques, to effectively address the challenges of large-scale missing data.
Large Language Model for MTSI
Large Language Models (LLMs) have catalyzed significant advancements in fields such as computer vision (CV) and natural language processing (NLP), and more recently in time series analysis Jin et al. (2024). LLMs, known for their exceptional generalization abilities, exhibit robust predictive performance, even when confronted with limited datasets. This characteristic is especially valuable in the context of MTSI. LLMs can adeptly mitigate these data gaps by leveraging multimodal knowledge, exemplified by their ability to incorporate additional textual information into analyses Jin et al. (2023), thus generating multimodal embeddings. Such a modeling paradigm not only enriches the imputation process by providing a more holistic understanding and representation of the data but also expands the horizons of MTSI. It enables the inclusion of varied data sources, thereby facilitating a more detailed and context-aware imputation. Exploring the integration of LLMs in MTSI represents a promising direction, with the potential to significantly enhance the efficacy and efficiency of handling missing data in multivariate time series data.

### v2

@@ 4.5 Pros and Cons
5 Large Model-based Methods
Large models aim to tackle three critical challenges in MTSI tasks: complex temporal dependencies across multiple scales, diverse missingness patterns, and the need for robust generalization with limited domain data. In this section, we examine how pre-trained foundation models (PFMs) and large language models (LLMs) approach these challenges through distinct but complementary paths.
5.1 Pre-Trained Foundation Models
PFMs leverage large-scale pretraining on diverse datasets to enhance generalization and adaptability Talukder et al. (2024). For example, by pretraining on a vast collection of multivariate time series, Timer Liu et al. (2024) learns a rich set of temporal representations, enabling it to generalize across different domains. It employs a self-supervised contrastive learning framework that refines its ability to reconstruct missing values while maintaining consistency in long-range dependencies. Building on the idea of large-scale pretraining, MOMENT Goswami et al. (2024) introduces the Time Series Pile and facilitates pretraining of high-capacity Transformer models using masked time series prediction tasks. On the other hand, NuwaTS Cheng et al. repurposes pre-trained language models for time series imputation, utilizing specialized embeddings and contrastive learning to handle various missing data patterns across domains. Meanwhile, Timemixer++ Wang et al. (2024a) explores an alternative to Transformer-based architectures by adopting token-mixing techniques. By mixing temporal and feature-wise representations, Timemixer++ effectively captures both short- and long-term dependencies while reducing computational overhead. These PFMs represent a significant step forward in MTSI, offering scalable, generalizable, and high-performance solutions that adapt to various missing data patterns across domains.
5.2 Large Language Models
LLMs have shown good capabilities in sequential modeling due to their autoregressive nature, which enables them to capture sequence dependencies through next-token prediction. Their extensive model parameters equip them with a strong learning capacity, making them suitable for handling MTSI tasks. For example, GPT4TS Zhou et al. (2023) optimizes GPT-2 for MTSI by introducing a key architectural adjustment—the freezing of the attention module. This design allows the model to focus on fine-tuning positional embeddings and layer normalization layers using a small number of samples. Such targeted fine-tuning significantly improves the model’s ability to reconstruct missing values with minimal data availability. Furthermore, LLM-TS Integrator Chen et al. (2024a) enhances LLM-based MTSI tasks by integrating statistical and deep learning methods within a hybrid framework. It introduces an adaptive retrieval mechanism to select relevant historical patterns and a self-correction module for iterative refinement. By leveraging retrieval-augmented generation, the model effectively handles irregular and sparse time-series data.
These large model approaches represent a fundamental shift in time series imputation: from treating missing values as isolated points to understanding them within a rich temporal and cross-variable context. While showing promising results, particularly in scenarios with complex missingness patterns, their adoption requires careful consideration of computational resources and domain-specific constraints.
@@ 6 Time Series Imputation Toolkits
On the MTSI task, there are existing libraries providing naive processing ways, statistical methods, machine learning imputation algorithms, and deep learning imputation neural networks for convenient usage.
@@
imputeTS Moritz and Bartz-Beielstein (2017), a library in R provides several naive approaches (e.g., mean values, last observation carried forward, etc.) and commonly-used imputation algorithms (e.g., linear interpolation, Kalman smoothing, and weighted moving average) but only for univariate time series. Another well-known R package, mice Van Buuren and Groothuis-Oudshoorn (2011), implements the method called multivariate imputation by chained equations to tackle missingness in data. Although it is not for time series specifically, it is widely used in practice for multivariate time-series imputation, especially in the field of statistics. Impyute [^2] https://github.com/eltonlaw/impyute and Autoimpute [^3] https://github.com/kearnz/autoimpute both offer naive imputation methods for cross-sectional data and time-series data. Impyute is only with simple approaches like the moving average window, and Autoimpute integrates parametric methods, for example, polynomial interpolation and spline interpolation. More recently, GluonTS Alexandrov et al. (2020), a generative machine-learning package for time series, provides some naive ways, such as dummy value imputation and casual mean value imputation, to handle missing values. In addition to simple and non-parametric methods, Sktime Löning et al. (2019) implements one more option that allows users to leverage integrated machine learning imputation algorithms to fit and predict missing values in the given data, though this works in a univariate way. ImputeBench Khayati et al. (2020) offers a collection of machine learning and deep learning-based imputation but lacks uniform programming languages.
When it comes to deep-learning imputation, PyPOTS Du (2023) is a toolbox focusing on modeling partially-observed time series end-to-end. It contains dozens of neural networks for tasks on incomplete time series, including 37 imputation models so far. Leveraging the toolkits from PyPOTS Ecosystem [^4] https://pypots.com/ecosystem that we developed, TSI-Bench Suite Du et al. (2024) provides a set of standard pipelines processing 172 public time-series datasets for benchmarking time-series imputation algorithms. Our comprehensive benchmark results from 34,804 experiments including 28 algorithms, 8 typical datasets from different domains, and diverse missingness patterns are presented in Du et al. (2024).
7 Future Direction
@@
Existing imputation algorithms predominantly operate under the MCAR or MAR assumptions. However, real-world missing data mechanisms are often more complex, with MNAR being prevalent across various domains. The non-ignorable nature of MNAR indicates a fundamental distributional shift between observed and true data. For example, in airflow signal analysis, the absence of high-value observations leads to saturated peaks, visibly skewing the observed data distribution compared to the true underlying one. This scenario illustrates how imputation methods may incur inductive bias in model parameter estimation and underperform in the presence of MNAR. Addressing missing data in MNAR contexts, distinct from MCAR and MAR, calls for innovative methodologies to achieve better performance.
@@
The primary objective of imputing missing values lies in enhancing downstream data analytics, particularly in scenarios with incomplete information. The prevalent approach is the “impute and predicttwo-stage paradigm, where missing value imputation is a part of data preprocessing and followed by task-specific downstream models (e.g. a classifier), either in tandem or sequentially. An alternative method is the “encode and predictend-to-end paradigm, encoding the incomplete data into a proper representation for multitask learning, including imputation and other tasks (e.g. classification and forecasting, etc.). Despite the optimal paradigm for partially-observed time series still remains an open area for future investigation, the latter end-to-end way turns out to be more promising especially when information embedded in the missing patterns is helpful to the downstream tasks Miyaguchi et al. (2021).
@@
While deep learning imputation algorithms have shown impressive performance, their computational cost often exceeds that of statistical and machine learning-based counterparts. In the era of burgeoning digital data, spurred by advancements in communication and IoT devices, we are witnessing an exponential increase in data generation. This surge, accompanied by the prevalence of incomplete datasets, poses significant challenges in training deep models effectively Wu et al. (2023b). Specifically, the high computational demands of existing deep imputation algorithms render them less feasible for large-scale datasets. Consequently, there is a growing need for scalable deep imputation solutions, leveraging parallel and distributed computing techniques, to effectively address the challenges of large-scale missing data.
Large Language Models for MTSI
While large models have shown promising results in time series imputation, several critical research directions remain unexplored. First, beyond current architectural innovations, the explicit incorporation of domain-specific temporal constraints and prior knowledge about missingness mechanisms into the pretraining process offers significant potential. Second, while the latest models have advanced temporal modeling, there remains room for fundamental innovations in processing irregular temporal patterns, particularly through more efficient and interpretable architectures. Third, the potential of multimodal learning deserves further investigation, where large models’ ability to process different data modalities could incorporate auxiliary information (such as textual descriptions or metadata) to achieve more accurate and contextually appropriate imputations. These directions, coupled with robust evaluation frameworks assessing uncertainty quantification and temporal consistency, could significantly advance time series imputation in critical applications.
8 Conclusion
This survey presents a systematic review of deep learning-based methods for multivariate time series imputation with a novel taxonomy to categorize predictive and generative methods and also discusses the large model for MTSI tasks. We provide a comprehensive architecture overview, highlighting their strengths, limitations, and applications. To advance this field, we identify key challenges, including handling MNAR missingness, integrating imputation with downstream tasks, and improving scalability. Future research should explore large-scale pre-trained models and multimodal learning to enhance robustness and real-world applicability.